OpenStudy (anonymous):

Equation(s) for the Vertical Asymptote(s): x-intercept in (x, y) form:

OpenStudy (anonymous):

g(x)=-log(x)+2

OpenStudy (jdoe0001):

usually the vertical asymptote for log(x) functions is 0, +2 will shift it over to the right

OpenStudy (anonymous):

I thought that +2 gave it a shift to the left.

OpenStudy (rajee_sam):

X-intercept can be found by putting y = 0 in the fn. 0 = -log(x) + 2 -2 = - log x 2 = log x ; x = 10² = 100 X-intercept is (100,0)

OpenStudy (rajee_sam):

+ 2 is a shift to top

OpenStudy (rajee_sam):

log (x -h) + k; h is the horizontal shift and k is the vertical shift. Since +2 is not with the x inside the log and since it is added after log(x) it is a vertical shift by 2 to the top since it is + ve

OpenStudy (rajee_sam):

y - intercept will be when x = 0 which is y = -1 + 2 = 1 Y-intercept is (0,1)